three coins are tossed let h be the random variable when head will occur.find the values of the random variable H


1
Expert's answer
2022-03-01T13:12:20-0500

There are "2^3=8" possible outcomes


"S=\{HHH, HHT, HTH, THH,"

"HTT, THT, TTH, TTT\}"

The possible values of the random variable "H" are "0, 1, 2, 3."

We will assume that the probability of getting heads and tails is the same:


"p = q =1\/2""\def\arraystretch{1.5}
 \begin{array}{c:c}
 Possible \ Outcomes & H \\ \hline
 HHH & 3 \\
 \hdashline
 HHT & 2 \\
 \hdashline
 HTH & 2 \\
 \hdashline
 THH & 2 \\
 \hdashline
 HTT & 1 \\
 \hdashline
 THT & 1 \\
 \hdashline
 TTH & 1 \\
 \hdashline
 TTT & 0 \\
 \hdashline 
\end{array}"


Construct the probability distribution of the random variable


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 h & 0 & 1 & 2 & 3 \\ \hline
 p(h) & 1/8 & 3/8 & 3/8 & 1/8
\end{array}"

or


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 h & 0 & 1 & 2 & 3 \\ \hline
 p(h) & 0.125 & 0.375 & 0.375 & 0.125 
\end{array}"